OpenStudy (anonymous):

Write the point slope form of the equation of the line described through (4,2), and parallel to y=-3/4x-5 I really don't get this can someone help :/

OpenStudy (baru):

can you identify the "slope" of the equation y=-3/4x-5?

OpenStudy (anonymous):

@kaiters1010 what happens to the 5 :?

hartnn (hartnn):

we are not changing the equation y=-3/4x-5, so nothing is gonna happen to 5. we are tying to find the equation of another line parallel to line y=-3/4x-5 so it will be in the form of y=-3/4x+c and we need to find the 'c' we can use the information that 4,2 lies on the line y=-3/4x+c so plug in x =4, y =2 in that 2 =-3/4* 4 +c ! once you get the 'c' , you get the equation

OpenStudy (anonymous):

and c is gotten like y=x+b correct?

hartnn (hartnn):

you mean y = mx+b ? here m is the slope , b is the y-intercept. we have slope = m= -3/4 here can you try to solve 2 =-3/4* 4 +b to get the value of the y-intercept 'b' ?

OpenStudy (anonymous):

b=-1

hartnn (hartnn):

how did you get that?? 2 = -(3/4)*4 + b 2 = -3 + b now add 3 on both sides :)

OpenStudy (anonymous):

Oh which it would be 1 not negative

hartnn (hartnn):

actually, lets do it step wise, 2 = -3 + b 2+3 = -3+3 +b 5 = 0 + b so b= 5 makes sense?

OpenStudy (anonymous):

Yes It did in the beginning I mixed up on the negative positive bit.

hartnn (hartnn):

so finally, we have the equation as y = -3/4 x + 5 :)

hartnn (hartnn):

ahh, just read the question, we need the equation in "point-slope" form! which is \(y = y_1 = m (x-x_1)\) and the point \(x_1, y_1 = 4,2 \) so the same equation would be \(y - 2 = -(3/4) (x- 4)\)

hartnn (hartnn):

*** \(\Large y \color{red}-y_1 = m (x-x_1)\)

OpenStudy (anonymous):

So x1 nd y1 is the one i got from the one that is parallel ?

OpenStudy (baru):

x1 and y1 are the co-ordinates that are given to you : (4,2)

OpenStudy (baru):

confused?

OpenStudy (baru):

(4,2) = (x1,y1)

OpenStudy (anonymous):

then what about y and x I'm getting really confused :/

OpenStudy (anonymous):

Would the solution be y−2=3/4(x−4)

hartnn (hartnn):

y−2=-3/4(x−4)

OpenStudy (anonymous):

I see and would that end of the question because it got very simple now

hartnn (hartnn):

yep, thats all we had to do, find the slope, which was -3/4 and plug in the point! x1 = 4, y1 =2

OpenStudy (anonymous):

So basically the point slope form would be y-2=-3/4(x-4

hartnn (hartnn):

yes

OpenStudy (anonymous):

Ok I got 3 more questions I'll try to do myself if I have any questions would it be ok to tag you :?

hartnn (hartnn):

sure :)